2. Find the capacity in litres of a conical vessel with
(i) radius 7 cm, slant heiiiht 25 cm
(ii) height 12 cm, slant height 13 cm.

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
(i) r = 7 cm, l = 25 cm. h = fl27 = (25)2 -(7r' cm = ../625- 49 cm = ../576 cm = 24 cm. .. Capacity = -i x x 7 x 7 x 24 em3 = 1232 cm3 = 1.232 l. (ii) h = 12 cm, l = 13 cm. r = l2>- h2 = (13)2 -(12)2 cm = ../169 -144 cm = cm = 5 cm. .. Capacity = -i x x 5 x 5 x 12 cm3 = 2200 cm3 = 11 litre. 7 35